OpenStudy (joehoffy):

What is the value of b in the equation 3b + 2(b − 1) = 8? 2 5 8 10

OpenStudy (joehoffy):

@anthonyym

Mehek (mehek14):

first distribute 2(b-1) = 2*b = ? 2*-1 = ?

Mehek (mehek14):

which one?

OpenStudy (joehoffy):

sorry lol -2

Mehek (mehek14):

that's 1 of them what abut 2*b?

OpenStudy (joehoffy):

thats the one idk

Mehek (mehek14):

2*b = 2b simple

Mehek (mehek14):

so 3b + 2b - 2 = 8 3b+2b=?

OpenStudy (joehoffy):

5b

Mehek (mehek14):

correct and add 8 + 2

OpenStudy (joehoffy):

10

Mehek (mehek14):

5b = 10 10/5 = ?

OpenStudy (joehoffy):

2

Mehek (mehek14):

that's your answer

OpenStudy (joehoffy):

THANKS!!

Mehek (mehek14):

yw :)